▼ Part A What is the period of a simple pendulum 70 cm long on the Earth? Express your answer using two significant figures. If there is no answer, enter none. SubmitP Previous Answers Request Answer X Incorrect: Try Again; 2 attempts remaining Part B at is the period of a simple pendulum 80 <cm long when it is in a freely falling elevator? Express your answer using two significant tigures. It there is no answer, enter none.
Answer #1

here,

A)

length of pedulum , l = 70 cm = 0.7 m

the period , T = 2*pi*sqrt(l/g)

T = 2*pi*sqrt(0.7/9.81)

T = 1.68 s

B)

for the pendulum in elevator

the net accelration is zero

the time period , T = 2*pi*sqrt(l/a)

T = infinite